Assume company ABC’s only product is priced at K20 per unit
and its variable costs amount to K15 per unit while fixed costs
are K3,600.
a. Compute the quantity required to achieve a profit of K2
per unit (5marks)
b. Define the term: Degree of Operating Leverage and how it
can be computed (5marks)
1
Expert's answer
2020-05-08T14:26:17-0400

a. Compute the quantity required to achieve a profit of K2 per unit

Let the quantity that is required to achieve the targeted profit per unit be X.

The total revenue Function will be "\text{Revenue}=20x"

The total cost to produce the given volume will be:

"\text{Cost}=3600+15x"

Therefore, if the profit per unit is 2 then the total profit will be "2x"

Thus the volume will be:

"2x=20x-(3600+15x)"

"2x-20x+15x=-3600"

"-3x=-3600"

"x=\dfrac{-3600}{-3}"


"X=1200 \text{Units}"

b. Define the term: Degree of Operating Leverage and how it can be computed

The degree of operating leverage is a measure that is used to measure the change in the level of the income of a company as a result of a change in the level of sales by a company.

The degree of operating can be computed using the following formula;

"\text{Degree of operating leverage}=\dfrac{\Delta \text{Operating income}}{\Delta \text{Sales}}"


Alternative formula that can be used is:

"\text{Degree of operating leverage}=\dfrac{\text{Contribution margin}}{\text{Operating income}}"